OpenStudy (stormswan):

Find 3x - y - 3z if x = -2, y = 1, and z = -2.

OpenStudy (stormswan):

im just checking if i got this one right....

OpenStudy (anonymous.):

Plug them in

OpenStudy (anonymous.):

3(-2) - 1 - 3(-2) = ?

OpenStudy (anonymous.):

-6-1-(-6) = ?

OpenStudy (stormswan):

-1....

OpenStudy (anonymous.):

-6-1-6 = ? :3

OpenStudy (stormswan):

-13

OpenStudy (anonymous.):

Correct :D

OpenStudy (stormswan):

well, thanks for your help :P

OpenStudy (anonymous.):

Your welcome :3